# SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver

### v2

@@ 3.2 Modifying models
Perturbations are given by de Putter et al. (2010). To implement this species in SymBoltz, write down all related variables, parameters, equations and initial conditions in one place: {codebox}
g, τ, k = M.g, M.τ, M.ka, ℋ, Φ, Ψ = g.a, g.ℋ, g.Φ, g.ΨD = Differential(τ)@parameters w₀ wₐ cₛ² Ω₀ ρ₀@variables ρ(τ) P(τ) w(τ) cₐ²(τ) δ(τ,k) θ(τ,k) σ(τ,k)eqs = [ w ~ w₀ + wₐ*(1-a) ρ₀ ~ 3*Ω₀ / (8*Num(π)) ρ ~ ρ₀ * a^(-3(1+w₀+wₐ)) * exp(-3wₐ*(1-a)) P ~ w * ρ cₐ² ~ w - 1/(3ℋ) * D(w)/(1+w) D(δ) ~ 3ℋ*(w-cₛ²)*δ - (1+w) * ( (1+9(ℋ/k)^2*(cₛ²-cₐ²))*θ - 3*D(Φ)) D(θ) ~ (3cₛ²-1)*ℋ*θ + k^2*cₛ²*δ/(1+w) + k^2*Ψ σ ~ 0]initialization_eqs = [ δ ~ -3//2 * (1+w) * Ψ θ ~ 1//2 * (k^2*τ) * Ψ]X = System(eqs, τ; initialization_eqs, name = :X)Everything is packed down in the $w_{0}w_{a}$ component X. Unicode code symbols are encouraged to maximize similarity with equations (e.g. Ω₀ over Omega_0), and SymBoltz will automatically render them in -compatible environments (e.g. notebooks).
A $w_{0}w_{a}\text{CDM}$ model is now built by replacing the cosmological constant species Λ in $\Lambda\text{CDM}$ by the $w_{0}w_{a}$ species X: {codebox}
@@
This is all the user has to do. The modification simply consists of writing down the equations verbatim. It could hardly be more compact and to the point. Under the hood, SymBoltz will:
  • create input hooks for setting the parameters $w_{0}$, $w_{a}$, $c_{s}^{2}$, $\Omega_{0}$,
  • move $(\tau)$-dependent functions to the background,
  • move $(\tau,k)$-dependent functions to the perturbations,
  • expand $\texttt{D(w)}=\mathrm{d}w/\mathrm{d}\tau$ and $\texttt{D(Φ)}=\mathrm{d}\Phi/\mathrm{d}\tau$ from $w$ and $\Phi$,
  • compute needed background variables in the perturbations,
  • spline $\rho(\tau)$ in the perturbations (if we use (7a) over (7b)),
  • source gravity with energy-momentum contributions,
  • assign $\Omega_{X0}=1-\sum_{s\neq{X}}\Omega_{s0}$ (if gravity is GR) by default,
  • eliminate common subexpressions like x=1+wx\!=\!1+w and y=1/xy\!=\!1/x,
  • generate ODE functions and state indices for $\delta^{\prime}$ and $\theta^{\prime}$,
  • generate sparse analytical Jacobian entries for $\delta^{\prime}$ and $\theta^{\prime}$,
  • interpolate and output any variable from the solution, both for unknowns (i.e. $\delta$ and $\theta$) and observeds (e.g. $w$ and $c_{a}^{2}$).
Most of these steps require manual intervention in a numerical code. For example, in CLASS one would have to read new parameters in input.c; declare new background and perturbation variables in background.h and perturbations.h; solve background equations, save desired output and source gravity in background.c; and recompute or look up background variables, solve perturbation equations, save desired output and source gravity in perturbations.c. The changes are scattered across the code, so it grows in complexity as more models are added. SymBoltz opens for a workflow that implements and analyzes a modified model with less effort all in one notebook, for example.
Of course, both codes already include a $w_{0}w_{a}$ species, but this is how they are implemented. We proceed with the $w_{0}w_{a}\text{CDM}$ model.
Figure 3: Matter and CMB ($T$, TE and $E$) power spectra computed by SymBoltz (SB; colored lines) compared to CLASS (CL; grey dashes) with relative errors $\text{rel.err.}=P_{k}^{\text{SB}}/P_{k}^{\text{CL}}-1$ and $\text{rel.err.}=C_{l}^{\text{SB}}/C_{l}^{\text{CL}}-1$ for the $\Lambda\text{CDM}$ model. CLASS uses the precision parameters in appendix B.
@@ 3.3 Computing power spectra
SymBoltz can compute the matter power spectrum $P(k)$ and the angular CMB power spectra $C_{l}^{\text{TT}}$, $C_{l}^{\text{TE}}$ and $C_{l}^{\text{EE}}$: {codebox}
ks = 10 .^ range(-1, 4, length = 100)Ps = spectrum_matter(prob, ks)ls = vcat([2,3,5,10], 20:20:2500)jl = SphericalBesselCache(ls)Cls = spectrum_cmb([:TT, :EE, :TE], prob, jl)The spectrum_matter and spectrum_cmb functions select a coarse $k$-grid to solve the perturbations for and interpolates to a finer $k$-grid. This is a common trick to integrate fewer $k$-modes. Precision parameters that affect the output are given with optional keyword arguments. Output spectra and their computation time with standard precision settings are shown in figs. 3 and 1.
It shows that SymBoltz and CLASS agree to around $0.1\%$ or better. The $P(k)$ agree to $0.1\%$ for all $k$ and $0.01\%$ for linear $k$. The $C_{l}$ agree to $0.1\%$ for most $l$, but slightly worse for cosmic variance-dominated $l$ and very large $l$ in $C_{l}^{\text{TE}}$. This is sufficient for most current data (e.g. Bolliet et al., 2024). It is similar to CAMB vs. CLASS with standard precision, although their agreement is pushed to $0.01\%$ with high precision (Lesgourgues, 2011b).
SymBoltz computes $P(k)$ quickly, but the extra steps needed for $C_{l}$ are not yet as optimized as they are in CLASS. A more thorough performance comparison for $P(k)$ is given in section 4.
@@ 3.4 Differentiable Fisher forecasting
Near a peak $\boldsymbol{p}_{0}$, where derivatives vanish, a log-likelihood $\log L$ of parameters $\boldsymbol{p}$ is approximated by the Taylor series
$\log L(\boldsymbol{p})\approx\log L(\boldsymbol{p}_{0})-\sum_{i,j}F_{ij}(\boldsymbol{p}_{0})(p_{i}-\bar{p}_{i})(p_{j}-\bar{p}_{j}),$ (8)
where $\boldsymbol{F}$ is the Fisher (information) matrix with elements
@@
Intuitively, $\boldsymbol{F}$ measures how sharp the peak is, or how sensitive $L$ is in different directions in parameter space. Fisher forecasting is powered by the Cramér-Rao bound $\big\lvert\smash{C_{ij}}\big\rvert\geq\big\lvert\smash{F^{-1}_{ij}}\big\rvert$ for the covariance $C_{ij}$ between parameters $p_{i}$ and $p_{j}$. It is an equality for a Gaussian, for which the likelihood (8) is exact. Thus, inverting $\boldsymbol{F}(\boldsymbol{p}_{0})$ gives the tightest parameter constraints one can hope for.
To demonstrate differentiable Fisher forecasting with SymBoltz, we make the best possible CMB (TT) measurement: one of $\bar{C}_{l}$ over the full sky with errors only due to cosmic variance
@@
The derivatives $\partial C_{l}/\partial p_{i}$ are usually found with error-prone finite differences by carefully tuning the step. SymBoltz avoids this problem altogether with automatic differentiation: {codebox}
vary = [ M.g.h, M.c.Ω₀, M.b.Ω₀, M.b.YHe, M.ν.Neff, M.h.m_eV, M.X.w₀, M.X.wₐ, M.I.ln_As1e10, M.I.ns,]genprob = parameter_updater(prob, vary)jl, ls = SphericalBesselCache(100:25:1000), 100:1000Cl(p) = spectrum_cmb(:TT, genprob(p), jl, ls)p₀ = map(par -> p[par], vary)dCl_dp = ForwardDiff.jacobian(Cl, p₀)Here vary orders the parameters to differentiate with respect to, and genprob generates a new problem with updated parameters p. This prepares a vector-to-vector function Cl(p) that is differentiated by ForwardDiff.jacobian to compute $\partial C_{l}/\partial p_{i}$ with dual numbers, as shown in fig. 4. Computing and inverting the Fisher matrix (12) forecasts the constraints in fig. 5. They are in excellent agreement with finite difference results from CLASS, but these required significant tuning of precision parameters and step sizes.
Differentiable Fisher forecasts have also been done with an Eisenstein-Hu transfer function fit by Campagne et al. (2023), and through an Einstein-Boltzmann solver by Hahn et al. (2024).
@@ 3.5 Differentiable MCMC sampling with supernova data
Finally, we show that SymBoltz can output differentiable results for gradient-based MCMC samplers, like the No-U-Turn Sampler (NUTS). It uses gradients to sample parameter space more efficiently than the ?blind? Metropolis-Hastings algorithm.
We predict apparent magnitudes $m(z)=M+\mu(z)$ of Type Ia supernovae at redshift $z$ and fit to 1048 Pantheon observations $(z_{i},m_{i})$ by Scolnic et al. (2018). Their standard absolute magnitude is $M\approx-19.3$, and $\mu(z)=5\lg(d_{L}(z)/10\text{ pc})$ is the distance modulus from the background-derived luminosity distance [^14] This expression is valid for any $\Omega_{k0}$ with complex $\operatorname{sinc}(x)=\sin(x)/x$, but can be split into branches for positive, negative and zero $\Omega_{k0}$.
@@
We define the likelihood $L$ by a multivariate normal distribution of the 1048 supernovae with (constant) covariance matrix $C$ given in detail by Scolnic et al. (2018). To compute $L$, we interface SymBoltz with the probabilistic programming framework Turing.jl (Fjelde et al., 2025), which allows a high-level description of the probabilistic model equivalent to the log-likelihood
@@
We use the NUTS sampler in Turing. The code for this example can be found in the paper’s notebook. For the flat $w_{0}\text{CDM}$ model ($\Omega_{k0}=0$, $w_{a}=0$, fixed $\Omega_{r0}$) it gives the constraints in fig. 6.
The differentiable computation is not yet fast enough for MCMCs with perturbation-derived spectra. The results here are not groundbreaking, but show that the differentiable pipeline is functional. This is a first step to differentiating the solver quickly, and we aim to accelerate this computation with future work.
Differentiable Boltzmann codes have yet to demonstrate this ability, but more advanced MCMCs with a differentiable background have been done by Campagne et al. (2023), for example.
Figure 6: Parameter inference on 1048 Type Ia supernovae from Pantheon data, using 5000 MCMC samples with the gradient-based NUTS sampler in Turing.jl and differentiable predictions from SymBoltz.
@@ 4 Performance comparison to CLASS
Figure 7: Work vs. precision for computing the matter power spectrum $P(k)$ with SymBoltz and CLASS. SymBoltz is approximation-free and run with several implicit ODE solvers. CLASS is run with approximations on/off, its implicit/explicit evolvers ndf15/rk and precision settings in appendix B, but tight coupling is always approximated. For each setup, $P(k)$ is computed by integrating perturbation $k$-modes with ODE solver tolerances $10^{-2}$-$10^{-9}$ that controls the adaptive stepping. For each tolerance, we record the best of 3 runtimes and the $L_{2}$-compressed error $\smash{\big(\sum_{i=1}^{N}\!\big(P(k_{i})/\bar{P}(k_{i})-1\big)^{2}/N\big)^{1/2}}$ relative to high-precision spectra $\bar{P}(k)$ from both codes with approximations off, tolerance $10^{-10}$ and FBDF/ndf15. Both codes solve a $w_{0}w_{a}\text{CDM}$ model with equal parameters, $N=114$ equal wavenumbers $k_{i}$, multipole cutoff $l_{\text{max}}=16$ for all species, default sampling of massive neutrino momenta and parallelize over $k$. Times are from a Linux laptop with Intel i7-12800H CPU.
@@
Figure 7 compares performance versus precision of the matter power spectrum $P(k)$ computed with SymBoltz and CLASS. This computation solves the background, thermodynamics and perturbations for several $k$ and reads off $P(k)$ at the final time. The work is dominated by the perturbations, so it is a relevant test of SymBoltz’ approximation-free treatment of perturbations.
@@
The comparison shows how approximations and the ODE solver affect integration of the Einstein-Boltzmann equations. Without approximations, SymBoltz with Rodas5P is more than $10\times$ faster than CLASS with its implicit ndf15 evolver at comparable precision. When CLASS uses approximations and explicit evolver, SymBoltz is as fast while remaining approximation-free.
@@
To make this comparison as fair as possible, we configured both codes to integrate identical $k$-modes (decided by CLASS) without $k$-interpolation and use the same multipole cutoff $l_{\text{max}}$ for all species. However, the codes sample massive neutrino momenta $q$ differently. This greatly impacts the number of perturbation equations (see section A.7). SymBoltz’ 4 default momenta gives 127 perturbation equations in total. CLASS’ default $\texttt{tol\_ncdm\_newtonian}=10^{-5}$ gives 11 momenta and 245 equations (before approximations). Increasing this tolerance to $10^{-3}$ gives 5 momentum bins, 143 equations and is around $245/143\approx 1.7$ times faster, but introduces errors beyond $1\%$ in $P(k)$. We therefore leave both codes with default $q$-sampling, but note that CLASS could be up to $2\times$ faster at unchanged precision with sparser sampling. This would shift its curves down by $\lg 2\approx 0.3$. Howlett et al. (2012) (in appendix A) also find a precise 3-4 point scheme in CAMB, and Lee et al. (2025) speed up CLASS using integral equations for massive neutrinos in the fork CLASSIER.
Despite lack of approximation schemes, SymBoltz is fast due to high-order implicit solvers like Rodas5P that take accurate and long steps, fast functions for $\boldsymbol{f}$ and the analytically generated $\boldsymbol{J}$, sparse matrix methods with a constant predetermined sparsity structure and efficient sampling of massive neutrino momenta.
@@
SymBoltz is currently slower than CAMB and CLASS for computing $C_{l}$. This is subject to more optimizations than $P(k)$, such as sampling of the source function $S(\tau,k)$, interpolation in $\tau$, $k$, and $l$, evaluation of spherical Bessel functions $j_{l}(x)$ and optimized line-of-sight quadrature. These are very well polished by CAMB and CLASS, and unmatched by SymBoltz at this time. A detailed performance analysis for $C_{l}$ is therefore not meaningful now (but see table 1). With more work, the $C_{l}$ optimizations can be added to SymBoltz. The purpose of the comparison made here is to show the viability of approximation-free perturbations, which was a major milestone in designing SymBoltz.

### v3

@@ 3.2 Modifying models
Perturbations are given by de Putter et al. (2010). To implement this species in SymBoltz, we write down all related variables, parameters, equations, and initial conditions in one place: {codebox}
g, τ, k = M.g, M.τ, M.ka, ℋ, Φ, Ψ = g.a, g.ℋ, g.Φ, g.ΨD = Differential(τ)@parameters w₀ wₐ cₛ² Ω₀ ρ₀=3Ω₀/8π@variables ρ(τ) P(τ) w(τ) cₐ²(τ) δ(τ,k) θ(τ,k) σ(τ,k)eqs = [ w ~ w₀ + wₐ*(1-a) ρ ~ ρ₀ * a^(-3(1+w₀+wₐ)) * exp(-3wₐ*(1-a)) P ~ w * ρ cₐ² ~ w - 1/(3ℋ) * D(w)/(1+w) D(δ) ~ 3ℋ*(w-cₛ²)*δ - (1+w) * ( (1+9(ℋ/k)^2*(cₛ²-cₐ²))*θ - 3*D(Φ)) D(θ) ~ (3cₛ²-1)*ℋ*θ + k^2*cₛ²*δ/(1+w) + k^2*Ψ σ ~ 0]initialization_eqs = [ δ ~ -3//2 * (1+w) * Ψ θ ~ 1//2 * (k^2*τ) * Ψ]X = System(eqs, τ; initialization_eqs, name=:X)
The last line packs everything into a $w_{0}w_{a}$ component X. Unicode symbols are encouraged to maximize similarity with equations (e.g., $\Omega_{0}$ over Omega_0), and SymBoltz automatically renders them in -compatible environments (e.g., notebooks). A full $w_{0}w_{a}\text{CDM}$ model is now built by replacing the cosmological constant species $\Lambda$ in $\Lambda\text{CDM}$ by the $w_{0}w_{a}$ species X: {codebox}
@@
The modification consists of writing down the equations compactly and directly. This is all the user must do, and under the hood SymBoltz will:
  • create input hooks for setting the parameters $w_{0}$, $w_{a}$, $c_{s}^{2}$, $\Omega_{0}$;
  • move $(\tau)$-dependent functions to the background;
  • move $(\tau,k)$-dependent functions to the perturbations;
  • expand $\texttt{D(w)}=\mathrm{d}w/\mathrm{d}\tau$ and $\texttt{D($\Phi$)}=\mathrm{d}\Phi/\mathrm{d}\tau$ from $w$ and $\Phi$;
  • compute needed background variables in the perturbations;
  • spline $\rho(\tau)$ in the perturbations (if we use Eq. (7a) over (7b));
  • source gravity with energy-momentum contributions;
  • assign $\Omega_{X0}=1-\sum_{s\neq{X}}\Omega_{s0}$ by default (if gravity is GR);
  • eliminate common subexpressions (e.g., $x\hskip-2.0pt=\hskip-2.0pt1\hskip-1.0pt+\hskip-1.0ptw$ and $y\hskip-2.0pt=\hskip-2.0pt1\hskip-1.0pt/\hskip-1.0ptx$);
  • generate ODE functions and state indices for $\delta^{\prime}$ and $\theta^{\prime}$;
  • generate analytical and sparse Jacobian entries for $\delta^{\prime}$ and $\theta^{\prime}$;
  • interpolate and output any variable from the solution, both for unknowns (i.e., $\delta$ and $\theta$) and observeds (e.g., $w$ and $c_{a}^{2}$).
More manual work is required in a purely numerical code. For example, in CLASS, we would have to read new parameters in input.c; declare new background and perturbation variables in background.h and perturbations.h; solve background equations, save desired output and source gravity in background.c; and recompute or look up background variables, solve perturbation equations, save desired output, and source gravity in perturbations.c. The changes are scattered across the code, so it grows in complexity as more models are added. SymBoltz enables a workflow that implements and analyzes a modified model with less effort all in one notebook, for example. Of course, both codes already include a $w_{0}w_{a}$ species, but this is how they are implemented. We proceed with the $w_{0}w_{a}\text{CDM}$ model to the next section.
Figure 3: Matter and CMB ($T$, $E$, and TE) power spectra computed by SymBoltz (colored lines) compared to CLASS (gray dashes) with relative errors $P(k)_{\text{SymBoltz}}/P(k)_{\text{CLASS}}-1$ and $C_{l,\text{SymBoltz}}/C_{l,\text{CLASS}}-1$ for the $w_{0}w_{a}\text{CDM}$ model. CLASS uses the precision parameters in Appendix B.
@@ 3.3 Computing power spectra
SymBoltz can compute the matter power spectrum $P(k)$, and the angular CMB power spectra $C_{l}^{\text{TT}}$, $C_{l}^{\text{EE}}$, and $C_{l}^{\text{TE}}$: {codebox}
ks = 10 .^ range(-1, 4, length = 100)Ps = spectrum_matter(prob, ks)ls = vcat([2,3,5,10], 20:20:2500)jl = SphericalBesselCache(ls)Cls = spectrum_cmb([:TT, :EE, :TE], prob, jl)The spectrum_matter and spectrum_cmb functions select a coarse $k$-grid to solve the perturbations for and interpolate to a finer $k$-grid. This is a common trick to integrate fewer $k$-modes. Precision parameters that affect the output are given with optional keyword arguments. Output spectra and their computation time with standard precision settings are shown in Fig. 3 and Table 1.
It shows that SymBoltz and CLASS agree to around $0.1\%$ or better. The $P(k)$ agree to $0.1\%$ for all $k$ and $0.01\%$ for linear $k$. The $C_{l}$ agree to $0.1\%$ for most $l$, although slightly worse for cosmic variance-dominated $l$ and very large $l$ in $C_{l}^{\text{TE}}$. This is sufficient for most current data (e.g., Bolliet et al., 2024). It is similar to CAMB versus CLASS with standard precision, although their agreement is pushed to $0.01\%$ with high precision (Lesgourgues, 2011b). SymBoltz computes $P(k)$ efficiently, but the extra steps needed for $C_{l}$ are not yet as optimized as they are in CLASS. Section 4 gives a more thorough performance comparison for $P(k)$.
@@ 3.4 Differentiable Fisher forecasting
Near a peak $\bar{\boldsymbol{p}}$, where derivatives vanish, a log-likelihood $\log L$ of parameters $\boldsymbol{p}$ is approximated by the Taylor series
$\log L(\boldsymbol{p})\approx\log L(\bar{\boldsymbol{p}})-\sum_{i,j}F_{ij}(\bar{\boldsymbol{p}})(p_{i}-\bar{p}_{i})(p_{j}-\bar{p}_{j}),$ (8)
where $\boldsymbol{F}$ is the Fisher information matrix with elements
@@
Intuitively, $\boldsymbol{F}$ measures how sharp the peak is or how sensitive $L$ is in different directions in parameter space. Fisher forecasting is powered by the Cramér–Rao bound $\big\lvert\smash{C_{ij}}\big\rvert\geq\big\lvert\smash{F^{-1}_{ij}}\big\rvert$ for the covariance $C_{ij}$ between parameters $p_{i}$ and $p_{j}$. It is an equality for a Gaussian, for which the likelihood expansion (8) is exact. Thus, inverting $\boldsymbol{F}(\bar{\boldsymbol{p}})$ gives the tightest possible parameter constraints.
To demonstrate differentiable Fisher forecasting with SymBoltz, we make the best possible CMB (TT) measurement of $\bar{C}_{l}$ over the full sky with errors only due to cosmic variance
@@
The derivatives $\partial C_{l}/\partial p_{i}$ are usually found with error-prone finite differences and careful step size tuning. SymBoltz avoids this problem altogether with automatic differentiation: {codebox}
vary = [ M.g.h, M.c.Ω₀, M.b.Ω₀, M.b.YHe, M.ν.Neff, M.h.m_eV, M.X.w₀, M.X.wₐ, M.I.ln_As1e10, M.I.ns,]genprob = parameter_updater(prob, vary)jl, ls = SphericalBesselCache(100:25:1000), 100:1000Cl(p) = spectrum_cmb(:TT, genprob(p), jl, ls)p₀ = map(par -> p[par], vary)dCl_dp = ForwardDiff.jacobian(Cl, p₀)Here, vary orders the parameters to differentiate with respect to, and genprob generates a new problem with updated parameters p. This prepares a vector-to-vector function Cl(p) that is differentiated by ForwardDiff.jacobian to compute $\partial C_{l}/\partial p_{i}$ with dual numbers, as shown in Fig. 4. Computing and inverting the Fisher matrix (12) forecasts the constraints in Fig. 5. They agree with finite difference results from CLASS, but these required significant tuning of precision parameters and step sizes. Differentiable Fisher forecasts have also been demonstrated with an Eisenstein–Hu transfer function fit by Campagne et al. (2023), and through the Einstein–Boltzmann solver DISCO-EB by Hahn et al. (2024).
@@ 3.5 Differentiable MCMC sampling with supernova data
Finally, SymBoltz can output differentiable results for gradient-based MCMC samplers, such as the No-U-Turn Sampler (NUTS). It uses gradients to sample parameter space more efficiently than the random walking Metropolis–Hastings algorithm.
We fit 1048 Pantheon observations $(z_{i},m_{i})$ of Type Ia supernovae by Scolnic et al. (2018) to predicted apparent magnitudes $m(z)=M+\mu(z)$ at redshift $z$. Their standard absolute magnitude is $M\approx-19.3$, and $\mu(z)=5\lg(d_{L}(z)/10\text{ pc})$ is the distance modulus of the background-derived luminosity distance [^17] This expression is valid for any $\Omega_{k0}$ with complex $\operatorname{sinc}(x)=\sin(x)/x$, but can be split into branches for positive, negative, and zero $\Omega_{k0}$.
@@
We define the likelihood $L$ by a multivariate normal distribution of the 1048 supernovae with covariance matrix $C$ given by Scolnic et al. (2018). To compute $L$, we use the probabilistic programming framework Turing.jl [^18] https://github.com/TuringLang/Turing.jl by Fjelde et al. (2025) and specify a probabilistic model equivalent to the log-likelihood
@@
We used the NUTS sampler in Turing. The code for this example can be found in the paper’s notebook. For the flat $w_{0}\text{CDM}$ model ($\Omega_{k0}=0$, $w_{a}=0$, fixed $\Omega_{r0}$), it gives the constraints in Fig. 6.
Differentiable computations are not yet fast enough for MCMCs with perturbation-derived spectra. These preliminary results show that the differentiable pipeline works, and we hope to speed it up with future work. Differentiable Boltzmann codes have yet to show use with gradient-based MCMC samplers, but Campagne et al. (2023) has done more advanced MCMCs with a differentiable background, for example.
Figure 6: Parameter inference on 1048 Type Ia supernovae from Pantheon data, using 5000 MCMC samples with the gradient-based NUTS sampler in Turing and differentiable predictions from SymBoltz.
@@ 4 Performance comparison to CLASS
Figure 7: Work versus precision for computing the matter power spectrum $P(k)$ with SymBoltz and CLASS. SymBoltz is approximation-free and run with several implicit ODE solvers. CLASS is run with approximations on/off, its implicit/explicit evolvers ndf15/rk, and precision settings in Appendix B, but tight coupling is always approximated. For each setup, $P(k)$ is computed by integrating perturbation $k$-modes with ODE solver tolerances $10^{-2}$-$10^{-9}$ that controls the adaptive stepping. For each tolerance, we record the best of 3 runtimes and the $L_{2}$-compressed error $\smash{\big(\sum_{i=1}^{N}\!\big(P(k_{i})/\bar{P}(k_{i})-1\big)^{2}/N\big)^{1/2}}$ relative to high-precision spectra $\bar{P}(k)$ from both codes with approximations off, tolerance $10^{-10}$, and FBDF/ndf15. Both codes solve a $w_{0}w_{a}\text{CDM}$ model with equal parameters, $N=114$ equal wavenumbers $k_{i}$, multipole cutoff $l_{\text{max}}=16$ for all species, default sampling of massive neutrino momenta, and parallelization over $k$. Times are from a Linux laptop with an Intel i7-12800H CPU.
@@
Figure 7 compares performance versus precision of the matter power spectrum $P(k)$ computed with SymBoltz and CLASS. This computation solves the background, thermodynamics, and perturbations for several $k$ and reads off $P(k)$ at the final time. It is a relevant test of SymBoltz’ approximation-free treatment of the perturbations, as they dominate the computational work.
@@
The comparison shows how approximations and the ODE solver affect the integration of the equations. Without approximations, SymBoltz with Rodas5P is more than $10\times$ faster than CLASS with its implicit ndf15 evolver at comparable precision. When CLASS uses approximations and its explicit rk evolver, SymBoltz is as fast while remaining approximation-free.
@@
To make this comparison as fair as possible, we configured both codes to integrate identical $k$-modes (decided by CLASS) without $k$-interpolation and use equal multipole cutoffs $l_{\text{max}}$ for all species. However, the codes sample massive neutrino momenta $q$ differently. This greatly impacts the number of perturbation equations (see Appendix A.7). SymBoltz’ 4 default momenta give 127 perturbation equations in total. CLASS’ default $\texttt{tol\_ncdm\_newtonian}=10^{-5}$ gives 11 momenta and 245 equations (before approximations). Increasing this tolerance to $10^{-3}$ gives 5 momenta and 143 equations. This runs around $245/143\approx 1.7$ times faster, but introduces errors beyond $1\%$ in $P(k)$. We therefore leave both codes with default $q$-sampling, but note that sparser sampling could make CLASS up to $2\times$ faster at unchanged precision and shift its curves down by $\lg 2\approx 0.3$. Howlett et al. (2012, in Appendix A) also found a precise three- to four-point scheme in CAMB, and Lee et al. (2025) sped up CLASS using integral equations for massive neutrinos.
@@
Despite the lack of approximation schemes, SymBoltz is fast thanks to high-order implicit solvers, fast generated functions for $\boldsymbol{f}$ and the analytical $\boldsymbol{J}$, sparse matrix methods with a constant precomputed sparsity pattern, and efficient sampling of massive neutrino momenta. However, SymBoltz currently computes $C_{l}$ slower than CAMB and CLASS. This computation is subject to more optimizations than $P(k)$, such as sampling of source functions $S(\tau,k)$, interpolation in $\tau$, $k$, and $l$, evaluation of spherical Bessel functions $j_{l}(x)$, and fast line-of-sight quadrature. These aspects are not yet as polished in SymBoltz as in CAMB and CLASS, so a detailed performance analysis for $C_{l}$ (beyond Table 1) is not meaningful now. Further work can add these optimizations to SymBoltz. The purpose of the comparison made here is to show the viability of approximation-free perturbations, which marks a milestone in the design of SymBoltz.